what is the slope of y=1/2x-3?
Answer by doukungfoo(195)

The equation of a line in slope intercept form is:
y = mx + b
m is the slope of the line
and b is the y intercept of the line

so the question asked is
What is the slope of y=1/2x-3
The equation of this line is already in slope intercept form so all we have
to do is identify which part is the slope.
we know that for y = mx + b
that m is the slope
so in y = 1/2x - 3
the slope is 1/2